OpenStudy (anonymous):

perfect square trinomial (x+2)^2 - (x-3)^2=35 ???

OpenStudy (llort):

I am assuming you want me to find x. It's 4...

OpenStudy (anonymous):

x^2+4x+4-(x^2-6x+9)=35 x^2+4x+4-x^2+6x-9=35 10x-5=35 10x=40 x=4

OpenStudy (anonymous):

x^2+4x+4-x^2+6x-9=35 10x-5=35 x=4

OpenStudy (anonymous):

i just saw it on youtube and the guy didnt finish the explanation so thats y im asking here